the formula for the time atraffic light remains yellow is t=1/8s +1, where t is the time in second and s is the speed limit in mph
a.solve the equation for s
b.what is the speed limit at a traffic ligth that remains yellow for 4.5 seconds

Answer by anjali_yamini(27)



t= 1/8s + 1
t=4.5
so, 4.5 = 1/8s + 1
or, 4.5 - 1 = 1/8s
or, 3.5 = 1/8s
or, 3.5 * 8 = s

28 = s

CHECK :
t= 1/8 * 28 + 1
t= (28 + 8) /8
t= 36 /8
t = 4.5